### v1

# Information Propagation in Predator-Prey Dynamics of Turbulent Plasma

Abstract
Magnetically confined fusion plasmas exhibit predator-prey-like cyclic oscillations through the self-regulating interaction between drift-wave turbulence and zonal flow. To elucidate the detailed mechanism and causality underlying this phenomenon, we construct a simple stochastic predator-prey model that incorporates intrinsic fluctuations and analyze its statistical properties from an information-theoretic perspective. We first show that the model exhibits persistent fluctuating cyclic oscillations called quasi-cycles due to amplification of intrinsic noise. This result suggests the possibility that the previously observed periodic oscillations in a toroidal plasma are not limit cycles but quasi-cycles, and that such quasi-cycles may be widely observed under various conditions. For this model, we further prove that information of zonal flow is propagated to turbulence. This information-theoretic analysis may provide a theoretical basis for regulating turbulence by controlling zonal flow.
pacs:
Valid PACS appear here
Introduction.
—Predator-prey dynamics can be observed ubiquitously in various systems such as ecosystems [1, 2, 3], genetic systems [4], fluids [5, 6, 7, 8], or even quantum systems [9]. Magnetically confined fusion plasmas also exhibit predator-prey-like cyclic oscillations in various situations, including the self-regulating process between zonal flow and drift-wave turbulence [10, 11, 12, 13], the intermediate phase (I-phase) during the low-to-high confinement (L-H) transition [14, 15, 16, 17, 18], the formation and collapse of internal transport barriers [19], and self-regulated oscillations of magnetic islands [20]. Among these, the interplay between zonal flow (predator) and drift-wave turbulence (prey) is believed to play a crucial role in suppressing anomalous heat and particle transport, and therefore, intensive attempts have been made to regulate turbulence by controlling zonal flow. Despite this importance, the detailed mechanism and causality underlying the cyclic oscillations have not been elucidated [12]. To deepen our understanding of the self-regulating process between zonal flow and turbulence, it is desirable to construct a minimal model that exhibits the cyclic oscillations and investigate the causality within the oscillations.
The main aim of this Letter is twofold. The first is to construct a simple model for the cyclic oscillations observed in plasma turbulence. Note that the standard two-variable predator-prey model proposed by Diamond et al. [21, 10], which incorporates the self-regulation mechanism between turbulence and zonal flow, does not exhibit a stable cyclic oscillation. To address this issue, we aim to construct a stochastic predator-prey model that incorporates the effects of intrinsic noise from the perspective of statistical physics. Since it has been pointed out in the field of ecology that it is essential to consider intrinsic fluctuations in predator-prey dynamics [22, 23], we expect that such a statistical physics approach provides useful insights into plasma turbulence. The second is to quantify the statistical causality between zonal flow and turbulence in this model through the lens of information theory. While several previous studies have attempted to quantify the statistical causality in plasma turbulence using information-theoretic quantities, such as (net) transfer entropy [24, 25] or information length [26, 27, 28, 29, 30, 31, 32], these quantities do not change sign under time reversal and thus can be nonzero even in an equilibrium state where all probability currents vanish. Instead, here we employ information flow, which is a pivotal concept within the framework of information thermodynamics [33, 34, 35, 36, 35, 37, 38] and has recently been applied to standard fluid turbulence [39, 40, 41].
The constructed model can be interpreted as a standard two-variable predator-prey model that takes into account intrinsic fluctuations. We show that the model exhibits persistent fluctuating cyclic oscillations, called quasi-cycles [22, 23], due to stochastic amplification. The quasi-cycles can be observed for a wide range of parameters, including parameter values consistent with experiments [14] and gyrokinetic simulations [42]. This result suggests the possibility that the previously observed periodic oscillations in a toroidal plasma are not limit cycles, but rather quasi-cycles. Our analysis of the resonance condition also implies that such quasi-cycles may be widely observed under various conditions.
For this model, we quantify the causality between zonal flow and turbulence using information flow. Specifically, we analytically show that the information on zonal flow is propagated to turbulence. This information propagation occurs as long as zonal flow and turbulence coexist, regardless of the presence or absence of quasi-cycles. Thus, our information-theoretic analysis may provide a theoretical foundation for the regulation of turbulence through the control of zonal flow.
Figure 1: (a) Schematic of the model. (b) Vector fields for parameter values $p=0.5,d=b=\gamma=1$ (left) and $p=2,d=b=\gamma=1$ (right). The color bar denotes the magnitude of the vector fields. The red markers indicate the stable fixed points, and the blue markers indicate the unstable fixed points. (c) The trajectory of the Langevin equation. The parameter values are $\epsilon=1/2,\beta=1,D^{W}=1/2$. The characteristic period is given by $2\pi/\omega_{c}\simeq 10$.
Setup.
[…]
$\partial_{t}p_{t}({\bm{n}})$ $=\sum_{\rho}\left[W_{\rho}({\bm{n}}|{\bm{n}}-{\bm{S}}_{\rho})p_{t}({\bm{n}}-{\bm{S}}_{\rho})\right$.
$\qquad\qquad\qquad\left.-W_{\rho}({\bm{n}}+{\bm{S}}_{\rho}|{\bm{n}})p_{t}({\bm{n}})\right].$ (1)
Here, $W_{\rho}({\bm{n}}+{\bm{S}}_{\rho}|{\bm{n}})$ denotes the transition rate from state $\bm{n}$} to state ${\bm{n}}+{\bm{S}}_{\rho}$ due to the reaction $\rho$, where ${\bm{S}}_{\rho}=(S^{Z}_{\rho},S^{W}_{\rho})$ denotes the stoichiometric vectors, which quantifies the change of $\bm{n}$} during a reaction of type $\rho$. For simplicity, we assume the mass action law [43]:
$W_{\rho}({\bm{n}}+{\bm{S}}_{\rho}|{\bm{n}})$ $=\Omega\rho\prod_{\alpha\in\{Z,W\}}\dfrac{1}{\Omega^{\nu^{\alpha}_{\rho}}}\dfrac{n^{\alpha}!}{(n^{\alpha}-\nu^{\alpha}_{\rho})!},$ (2)
[…]
System size expansion.
—Since we are interested in the collective behavior of the system, we consider coarse-grained descriptions by taking the infinite volume limit $\Omega\rightarrow\infty$. Here, we note that the two limits $t\rightarrow\infty$ and $\Omega\rightarrow\infty$ do not commute in this model, which is known as Keizer’s paradox [44, 45]. Indeed, if we take the limit $t\rightarrow\infty$ first, then the probability distribution converges to a unique stationary distribution $p_{\mathrm{ss}}({\bm{n}})=\delta_{{\bm{n}},{\bm{0}}}$, which implies that eventually there will be no zonal flows and turbulences. Below, we focus on the opposite case, where we first take $\Omega\rightarrow\infty$ and then $t\rightarrow\infty$. The coarse-grained dynamics in the limit $\Omega\rightarrow\infty$ can be systematically obtained by applying van Kampen’s system size expansion [46]. To separate the fluctuating part from the average part, we introduce a new stochastic variable ${\bm{r}}_{t}=(z_{t},w_{t})\in\mathbb{R}^{2}$ as ${\bm{n}}_{t}/\Omega={\bm{c}}_{t}+{\bm{r}}_{t}/\Omega^{1/2}$, where ${\bm{c}}_{t}=(c^{Z}_{t},c^{W}_{t})\in\mathbb{R}^{2}_{\geq 0}$ denotes the mean energy density to be determined. By substituting this expression into Eq. (1) and expanding in terms of $\Omega^{-1/2}$, we find that the leading order $O(\Omega^{1/2})$ yields the time evolution equation for the average part ${\bm{c}}_{t}$, which can be regarded as a coarse-grained dynamics of Eq. (1):
$\dfrac{d}{dt}c^{Z}_{t}$ $=pc^{Z}_{t}c^{W}_{t}-dc^{Z}_{t},$ (3)
$\dfrac{d}{dt}c^{W}_{t}$ $=bc^{W}_{t}-\gamma(c^{W}_{t})^{2}-pc^{Z}_{t}c^{W}_{t}.$ (4)
See Ref. [47] for the detailed derivation. This equation corresponds to the standard two-variable predator-prey model proposed by Diamond et al. [21, 10], which has the same form as the Lotka–Volterra equation. There are two regimes in this model (Fig. 1(b)). When $bp<d\gamma$, the stable fixed point is $(c^{Z}_{\mathrm{ss}},c^{W}_{\mathrm{ss}})=(0,b/\gamma)$, which represents a state without zonal flows. In contrast, when $bp>d\gamma$, the stable fixed point is $((bp-d\gamma)/p^{2},d/p)$, which represents a coexistence state. Importantly, this equation does not exhibit limit cycles, contrary to the naive expectation of the existence of predator-prey oscillations. While cyclic oscillations emerge when $\gamma=0$, these cycles are not limit cycles but structurally unstable.
Note that the coarse-grained description [Eqs. (3) and (4)] ignores intrinsic noise that is inevitably present in the original individual-level description (1). For a realistic situation where $\Omega$ is large but finite, it becomes essential to consider the intrinsic noise. To investigate the effect of the intrinsic noise on the coexistence state, we focus on the regime $bp>d\gamma$ and proceed to the subleading order $O(\Omega^{0})$ of the system size expansion by substituting $(c^{Z}_{\mathrm{ss}},c^{W}_{\mathrm{ss}})=((bp-d\gamma)/p^{2},d/p)$. The resulting equation describes the time evolution of the fluctuating part ${\bm{r}}_{t}$ around ${\bm{c}}_{t}={\bm{c}}_{\mathrm{ss}}$ [47]:
$\dfrac{d}{d\tau}z_{\tau}$ $=\epsilon w_{\tau}+\zeta^{Z}_{\tau},$ (5)
$\dfrac{d}{d\tau}w_{\tau}$ $=-\beta z_{\tau}-(1-\epsilon)w_{\tau}+\zeta^{W}_{\tau},$ (6)
where $\tau:=bt$ denotes the dimensionless time, $\beta:=d/b$ denotes the dimensionless damping rate, and $\epsilon:=1-d\gamma/bp\in(0,1)$ denotes the ratio of time scales for $z$ and $w$. The terms $\zeta^{Z}_{\tau}$ and $\zeta^{W}_{\tau}$ denote the zero-mean white Gaussian noise that satisfies $\langle\zeta^{Z}_{\tau}\zeta^{Z}_{\tau^{\prime}}\rangle=2D^{Z}\delta(\tau-\tau^{\prime})$, $\langle\zeta^{Z}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=-D^{Z}\delta(\tau-\tau^{\prime})$, and $\langle\zeta^{W}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=2D^{W}\delta(\tau-\tau^{\prime})$, where $D^{Z}:=\epsilon c^{W}_{\mathrm{ss}}=\epsilon d/p$ and $D^{W}:=c^{W}_{\mathrm{ss}}=d/p$. Note that this linear Langevin equation has a form in which noise is added to the linearized Lotka–Volterra equation around the coexistence fixed point. We emphasize that the noise intensities $D^{Z}$ and $D^{W}$ are uniquely determined from the master equation (1). In other words, the noise driving the system is not artificially added but rather arises from intrinsic stochasticity. This property marks a stark difference from the stochastic models used in Refs. [26, 27, 28, 29, 30, 31, 32], where noise is artificially added to deterministic predator-prey models. The Langevin equations (5) and (6) are equivalent to the Fokker–Planck equation $\partial_{\tau}p_{\tau}({\bm{r}})=-\partial_{z}J^{Z}_{\tau}({\bm{r}})-\partial_{w}J^{W}_{\tau}({\bm{r}})$, where $p_{\tau}({\bm{r}})$ denotes the probability density of finding ${\bm{r}}_{\tau}={\bm{r}}$ and $J^{Z}_{\tau}({\bm{r}})$ and $J^{W}_{\tau}({\bm{r}})$ denote the probability currents [47].
Quasi-cycles induced by stochastic amplification.
—Although the coarse-grained deterministic description [Eqs.(3) and (4)] cannot predict stable cycles, the stochastic description [Eqs. (5) and (6)] predicts persistent cyclic oscillations called quasi-cycles. Figure 1(c) shows an example of such persistent fluctuating oscillations observed in the linear Langevin equation. Notably, the fluctuations of zonal flow follow those of turbulence with a phase lag $\sim\pi/2$, which are similar to the predator-prey-like oscillations observed in a toroidal plasma [14, 15, 16, 17, 12]. This oscillatory behavior occurs when the coexistence state is a stable spiral fixed point in the coarse-grained deterministic description. More precisely, from the calculation of the power spectral density [47], we can show that the resonant stochastic amplification occurs for a wide range of parameters, where the characteristic resonant frequency $\omega_{c}:=\sqrt{\epsilon\beta-(1-\epsilon)^{2}/2}$ takes a real value (see also Fig. 2(b)). Because the white noise in the Langevin equation covers all frequencies, the corresponding cycle with the resonant frequency $\omega_{c}$ is amplified. This mechanism is called stochastic amplification to distinguish it from stochastic resonance [22].
The quasi-cycles differ from limit cycles in that they are stochastic oscillations where the different frequencies near $\omega_{c}$ are all excited [23, 48]. Furthermore, because quasi-cycles can occur as long as the system has a stable spiral fixed point and noise, they may be widely observed in plasma turbulence under various conditions. Although we have emphasized that our model is a qualitative description, here we present a set of parameter values consistent with experiments and simulations for reference; By comparing with the gyrokinetic simulations [42], they can be estimated as $\epsilon\sim 1$ and $\beta\sim 10^{-4}$, which satisfies the resonance condition, and the corresponding cycle frequency reads $\omega_{c}\sim 10^{-2}$ in dimensionless units.
Information propagation between zonal flow and turbulence.
—The presence of quasi-cycles suggests that intrinsic fluctuations are essential for understanding the dynamics of zonal flow and turbulence. Therefore, we scrutinize the statistical causality underlying the intrinsic fluctuations from an information-theoretic viewpoint.
We first introduce the mutual information [49], which quantifies the mutual dependence between zonal flow and turbulence:
$I[Z\colon\!W]:=\int dzdwp_{\tau}(z,w)\ln\dfrac{p_{\tau}(z,w)}{p^{Z}_{\tau}(z)p^{W}_{\tau}(w)}\geq 0,$ (7)
where $p_{\tau}(z,w)$ denotes the joint probability density, and $p^{Z}_{\tau}(z)$ and $p^{W}_{\tau}(w)$ denote the marginal distributions. The mutual information is nonnegative and equal to zero if and only if zonal flow and turbulence are statistically independent.
Figure 2: (a) Schematic of information flow $\dot{I}^{Z\rightarrow W}_{\lambda}$. (b) $(\epsilon,\beta)$-dependence of $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$. The dashed line denotes the resonance condition $2\epsilon\beta-(1-\epsilon)^{2}=0$. The stochastic amplification occurs in the parameter region above this line.
The decomposition of the time derivative of the mutual information yields information flow. Specifically, the information flow from zonal flow to turbulence is defined as (see Fig. 2(a))
$\dot{I}^{Z\rightarrow W}_{\lambda}:=\lim_{\Delta\tau\rightarrow 0^{+}}\dfrac{I[Z_{\tau+\lambda\Delta\tau}\colon\!W_{\tau+\Delta\tau}]-I[Z_{\tau+\lambda\Delta\tau}\colon\!W_{\tau}]}{\Delta\tau}$ (8)
with $\lambda\in[0,1]$. We can similarly define the information flow from turbulence to zonal flow $\dot{I}^{W\rightarrow Z}_{\lambda}$, which is related to $\dot{I}^{Z\rightarrow W}_{\lambda}$ as $d_{\tau}I[Z\colon\!W]=\dot{I}^{Z\rightarrow W}_{\lambda}+\dot{I}^{W\rightarrow Z}_{1-\lambda}$. Hence, we have $\dot{I}^{Z\rightarrow W}_{\lambda}=-\dot{I}^{W\rightarrow Z}_{1-\lambda}$ in the steady state. Here, the information flow $\dot{I}^{Z\rightarrow W}_{\lambda}$ depends on $\lambda$ because the Langevin equations (5) and (6) do not satisfy the bipartite condition. That is, the noises in Eqs. (5) and (6) are not independent, but correlated as $\langle\zeta^{Z}_{\tau}\zeta^{W}_{\tau^{\prime}}\rangle=-D^{Z}\delta(\tau-\tau^{\prime})$. We remark that, in the context of information thermodynamics [35, 37, 38], it is common to consider bipartite systems, where information flow does not depend on $\lambda$, and non-bipartite systems have rarely been studied [36].
The information flow $\dot{I}^{Z\rightarrow W}_{\lambda}$ quantifies the rate at which turbulence gains information about zonal flow. In particular, its sign indicates the direction in which the information is transmitted: if $\dot{I}^{Z\rightarrow W}_{\lambda}>0$ for all $\lambda$, the information of zonal flow is transferred to turbulence, and vice versa. Here, we note that the non-bipartite structure allows the case where the sign of $\dot{I}^{Z\rightarrow W}_{\lambda}$ varies depending on $\lambda$, which makes it difficult to determine the direction of information transfer [36]. In that case, $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$ plays a special role because it is odd under time reversal and vanishes in the equilibrium state. Fortunately, as will be shown below, this problem is irrelevant to our model.
From Eq. (8), we can express the information flow in terms of the probability current $J^{W}_{\tau}({\bm{r}})$ as
$\dot{I}^{Z\rightarrow W}_{\lambda}$ $=\int dzdw\left[J^{W}_{\tau}({\bm{r}})+(2\lambda-1)\dfrac{D^{Z}}{2}\dfrac{\partial}{\partial z}p_{\tau}({\bm{r}})\right]$
$\qquad\times\dfrac{\partial}{\partial w}\ln\dfrac{p_{\tau}({\bm{r}})}{p_{\tau}(z)p_{\tau}(w)},$ (9)
where $p_{\tau}(z)$ and $p_{\tau}(w)$ are marginal distributions for zonal flow and turbulence, respectively. By noting that the steady-state distribution $p_{\mathrm{ss}}({\bm{r}})$ for the linear Langevin equations (5) and (6) is Gaussian with covariance matrix $\Sigma:=\langle({\bm{r}}-\langle{\bm{r}}\rangle)({\bm{r}}-\langle{\bm{r}}\rangle)^{\top}\rangle$, the information flow in the steady state can be calculated as
$\dot{I}^{Z\rightarrow W}_{\lambda}$ $=\dfrac{D^{Z}}{\det\Sigma}\left[\Sigma^{WW}+(1-\lambda)\Sigma^{ZW}\right]$
$>0\quad\text{for all}\quad\lambda\in[0,1],$ (10)
where $\Sigma^{ZW}=\langle(z-\langle z\rangle)(w-\langle w\rangle)\rangle$ and $\Sigma^{WW}=\langle(w-\langle w\rangle)^{2}\rangle$. See Ref. [47] for the detailed derivation of Eqs. (9) and (10). The inequality (10) states that the information about fluctuations of zonal flow is propagated to turbulence. In other words, the turbulence is “learning” about the fluctuating behavior of the zonal flow. This result is somewhat counterintuitive because the system exhibits persistent predator-prey cycles, where the fluctuations of zonal flow follow those of turbulence with a time lag. Figure 2(b) shows the $(\epsilon,\beta)$-dependence of $\dot{I}^{Z\rightarrow W}_{\lambda=1/2}$. Interestingly, Fig. 2(b) suggests that the behavior of the information flow does not depend significantly on whether stochastic amplification occurs or not. In other words, even when the system does not exhibit the fluctuating predator-prey oscillation, there is a statistical causality from zonal flow to turbulence. This information-theoretic analysis may provide a theoretical foundation for the regulation of turbulence through the control of zonal flow. We finally remark that the inferred causality from the information flow is consistent with that from the cross-correlation function, although interpreting the latter is less straightforward [47].
Concluding remarks.
—We have proposed a simple stochastic predator-prey model for plasma turbulence that exhibits quasi-cycles induced by stochastic amplification of intrinsic noise. For this model, we have quantified the causality between zonal flow and turbulence by showing that the information of zonal flow is propagated to turbulence.
Our model illustrates the significance of considering the effects of intrinsic noise in understanding the predator-prey-like oscillations observed in plasma turbulence. Notably, our results suggest the possibility that the observed cyclic oscillations in a toroidal plasma are quasi-cycles rather than limit cycles. Using the methods proposed in Ref. [48], it would be possible to distinguish quasi-cycles and limit cycles both in experiments and simulations. Further studies are needed to verify whether both quasi-cycles and information transfer from zonal flow to turbulence can be observed in experiments and simulations of more realistic models, such as the Hasegawa–Wakatani equation [50].
Acknowledgements.
We thank S. Maeyama and M. Nakata for fruitful discussions. T.T. was supported by JSPS KAKENHI Grant Number JP25K17315 and JST PRESTO Grant Number JPMJPR23O6, Japan. M.S. was supported by JSPS KAKENHI Grant Number JP25K00986, Japan.